# RULSurv: A probabilistic survival-based method for early censoring-aware prediction of remaining useful life in ball bearings

### v2

@@ Abstract
Censored data refers to situations where the full information about a particular event or process is only partially known. In survival analysis, censoring plays an important role, as ignoring such observations can bias the model parameters and overestimate the probability of when the event is likely to occur. There has been a renewed interest in using data-driven methods to predict the remaining useful life (RUL) of ball bearings for predictive maintenance. However, few studies have explicitly addressed the challenge of handling censored data. To address this issue, we introduce a novel and flexible method for early fault detection using Kullback-Leibler (KL) divergence and RUL estimation using survival analysis that naturally supports censored data. We demonstrate our approach in the XJTU-SY dataset using a 5-fold cross-validation across three different operating conditions. When predicting the time to failure for bearings under the highest load (C1, 12.0 kN and 2100 RPM) with 25% random censoring, our approach achieves a mean absolute error (MAE) of 14.7 minutes (95% CI 13.6-15.8) using a linear CoxPH model, and an MAE of 12.6 minutes (95% CI 11.8-13.4) using a nonlinear Random Survival Forests model, compared to an MAE of 18.5 minutes (95% 17.4-19.6) using a linear LASSO model that does not support censoring. Moreover, our approach achieves a mean cumulative relative accuracy (CRA) of 0.7586 over 5 bearings under the highest load, which improves over several state-of-the-art baselines. Our work highlights the importance of considering censored observations as part of the model design when building predictive models for early fault detection and RUL estimation.
@@ 1 Introduction
However, despite numerous advances in model development and increases in predictive performances, one remaining challenge is how to handle censored (aka. incomplete) data, where the event of interest is only partially observed – i.e., for some instances, we only have a lower bound on the actual time of the event. Censored or incomplete observations are a prevalent phenomenon given the rarity of failures or breakdowns in ball bearings [15], but so far, only few studies have attempted to model or address the consequences of censoring [68, 67, 22, 75] and the literature on censored data is limited. We can choose to ignore censored observations and analyze only the data where we know the true outcome, but this would lead to a loss of information and introduce an estimation bias in the model parameters [56]. If we want to accurately estimate when a bearing is likely to break down based on historical data and current sensor readings, we should include information from censored data samples in the model. Moreover, since bearing failure is generally a rare event, in the sense that the number of data samples where the bearing operates without problems far exceeds the number of samples with failure indications [15], it is important to make use of the data we have at our disposal, including censored and uncensored observations.
@@ 2.2 RUL prediction
[75] proposed a time-dependent survival neural network (TSNN), which estimates the failure risk of a mechanical component by performing multiple classification steps in succession. The output of TSNN is an RUL-specific probability, which is monotonically decreasing. Their approach incorporates temporal dependencies between measurement points, and inherently supports censoring, but the method does not predict the time to failure, rather the probability of remaining failure-free by some time in the future. Additionally, the TSNN model requires a number of observational data points at prediction time. However, observational data require determining features beyond the observed time frame. As a result, time-variant predictions are difficult to make and survival bias may be introduced [79].
@@
[71] proposed a novel health indicator framework for automatic RUL prediction based on a multi-head attention architecture, combining information from damaged and healthy bearings to improve prediction accuracy. Their method achieves a good prediction error compared to other works, but the estimated RUL is not monotonically decreasing, and the authors make the assumption of a linear correlation between the RUL and a bearing’s lifetime. Assuming a linear RUL oversimplifies the bearing degradation process, as the relationship between the RUL and the number of life cycles is often nonlinear [28, Ch. 2]. However, bearings may not degrade at a constant speed, as their health can depend on various factors, e.g., operating condition, the type of maintenance performed, or other external factors [77].
@@
[37] proposed a hybrid deep learning approach for RUL prediction that utilizes non-full lifecycle (i.e., censored) data. The authors combine self-supervised and supervised learning to address data scarcity and noise challenges in RUL prediction tasks. Using a Contrastive Predictive Coding (CPC) encoder and a Transformer decoder, the model leverages non-full lifecycle data for pretraining and finetuning with full lifecycle data. Their experiments show good performance in a separate dataset in terms of the root mean squared error (RMSE). Although censored data are used in the pretraining phase, the prediction task in this case is simply predicting the next feature value $x_{t+1}$ given $x_{t}$, not the time to event. In other words, the censored data is not used to predict the probability of future event, but rather to improved the learned feature presentation. Moreover, the authors never evaluate their method with censoring in the testing phase and the proposed RMSE assumes that all samples have observed failure times and the predicted RUL is not monotonically decreasing.
@@ 3.2 Defect evolution and failure monitoring
The early stages of a bearing defect can be effectively detected from weak vibration signals through the enveloped Fourier transform [48], as this method can detect amplitude-modulated signals. By selecting an impulsive frequency band, often containing one or more of the component’s eigenfrequencies and applying the Hilbert transform, the low-frequency envelope or carrier frequency can be found [47, Ch. 3]. The carrier frequency corresponds to one of the four characteristic bearing frequencies that are bound to the bearing geometry, the number of rolling elements, and the rotation speed.
@@ 3.3 Elements of survival analysis
We define a survival problem by a sequence of observations represented as triplets ($\bm{x}_{i}$, $t_{i}$, $\delta_{i}$), where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a vector with $d$ features, $t_{i}$ denotes the time to event or censoring, and $\delta_{i}$ is a binary event indicator. Moreover, let $c_{i}$ denote the censoring time and $e_{i}$ denote the event time for the $i$th observation, thus $t_{i}=e_{i}$ if $\delta_{i}=1$ or $t_{i}=c_{i}$ if $\delta_{i}=0$. We consider the survival time as discrete and limit the time horizon to a finite duration, denoted $(0,...,T_{max})$, where $T_{max}$ represents a predefined maximum time point (e.g., 1 year). A survival model then estimates the probability that the event of interest occurs at time $T$ later than $t$, that is, the survival probability $S\left(t\right)=\text{P}\left(T>t\right)=1-\text{P}\left(t\leq T\right)$. To estimate the survival probability, we can use the hazard function:
@@
which represents the failure rate at an instant after time $t$, assuming survival beyond that time [18, Ch. 11]. The hazard function is related to the survival function through $h\left(t\right)=f\left(t\right)/S\left(t\right)$, where $f\left(t\right)$ is the probability density associated with $T$. Formally, $f\left(t\right):=\lim_{\Delta t\rightarrow 0}\text{P}\left(t<T\leq t+\Delta t\right)/\Delta t$, that is, the instantaneous failure rate at time $t$. Under this definition, the function $h\left(t\right)$ is the probability density of $T$ conditioned on $T>t$, where the higher the value of $h\left(t\right)$, the higher the probability of failure. The functions $S\left(t\right)$, $h\left(t\right)$, and $f\left(t\right)$, are different but related ways to describe the probability distribution of $T$.
@@ 3.4 Elements of uncertainty estimation
Many machine learning methods for RUL prediction are indeed nonprobabilistic; they can only provide a point estimate of the RUL and disregard the inherent uncertainty associated with the prediction. The Bayesian framework offers a theoretical approach to model uncertainty when estimating the model parameters (epistemic) and when making the actual prediction (aleatoric) [39]. This enables us to assess confidence in the obtained predictions, which further encourages the adoption of the model as a decision support tool. [42] introduced Bayesian Neural Networks (BNNs) for bearing health prognostics with uncertainty estimation to predict the RUL of turbofan engines. In a BNN, the weights are treated as a set of random variables $\bm{\theta}\sim\text{P}\left(\bm{\theta}\right)$. In this context, let $\mathcal{D}$ denote the data, $\text{P}\left(\mathcal{D}|\bm{\theta}\right)$ the likelihood and $\text{P}\left(\bm{\theta}\right)$ the prior distribution over the parameters of interest $\bm{\theta}$. Bayesian inference computes the posterior distribution $\text{P}\left(\bm{\theta}|\mathcal{D}\right)=\text{P}\left(\bm{\theta}\right)\text{P}\left(\mathcal{D}|\bm{\theta}\right)/\text{P}\left(\bm{\theta}\right)$, and then obtains the predictive distribution of unobserved data conditioned on $\mathcal{D}$ by marginalizing over the parameter space $\Theta$, i.e.$\text{P}\left(\bm{x}_{\text{new}}|\mathcal{D}\right)=\int_{\Theta}\text{P}\left(\bm{x}_{\text{new}}|\mathcal{D},\bm{\theta}\right)\text{P}\left(\bm{\theta}|\mathcal{D}\right)\text{d}\bm{\theta}$. However, the posterior cannot be directly computed due to $\text{P}\left(\mathcal{D}\right)$ being intractable, and sampling methods do not scale well in high dimensions. To this end, variational inference (VI) has been proposed to approximate the posterior with a tractable parametric distribution, $q_{\bm{\zeta}}\left(\bm{\theta}\right)$. This $q_{\bm{\zeta}}\left(\bm{\theta}\right)$ is chosen within a class of tractable parametric distributions, $\mathcal{Q}$, by minimizing the KL divergence from $q_{\bm{\zeta}}\left(\bm{\theta}\right)$ to $\text{P}\left(\bm{\theta}|\mathcal{D}\right)$. In many modern neural networks, Monte Carlo Dropout (MCD) or another timely technique is used to approximate the VI solution [17][34] proposed BNNs as a tool for uncertainty estimation in deep survival networks using MCD and showed a performance advantage in adopting the Bayesian framework, especially in small datasets, over nonprobabilistic models. The intrinsic probabilistic dimension of BNNs naturally allows us to make uncertainty estimates, estimate predictive errors, and plot credible intervals [45]. This, in turn, can justify the level of trust and confidence that practitioners have in the prediction. Nonprobabilistic models cannot give these advantages.
@@
Figure 2: Outline of the proposed method. Historical data for 5 ball bearings form a bearing dataset $\mathcal{D}$ (see Section 4.1). We first perform feature extraction of all the bearings in the time-domain (see Section 4.2). Then, to train a survival model for RUL prediction, we designate training and test bearings and perform event detection in the frequency-domain of the training bearings (see Section 4.3). Afterwards, we convert the temporal training dataset to a supervised learning dataset by computing a rolling average (see Section 4.4. We then apply random, independent censoring using the proposed censoring algorithm (see Section 4.5), so that the event of interest is only observed for a portion of the instances. This processed dataset $\tilde{\mathcal{D}}$ thus contains $N$ instances (rows) with several time-domain features (e.g., entropy, kurtosis), and the time to event (filled dot) or censoring (hollow dot). We use this dataset (features and event information) to train a survival model $\mathcal{M}$ (see Section 4.6) that can estimate the individual survival distribution (ISD) of the failure event, for a new test bearing $\bm{x}_{i}$, denoted as $\hat{s}_{i}$. This ISD give the probability that failure (i.e., the onset of degradation) occurs after $t$ minutes post startup, for all $t>0$. It can also be used to estimate the time to event for this $\bm{x}_{i}$, for example, when the survival (event) curve intersects the dashed horizontal line at 50%, which is referred to as the predicted median survival time.
@@ 4.1 The dataset
We use the publicly available Xi’an Jiaotong University (XJTU-SY) [63] dataset to evaluate the performance of our proposed method. The dataset contains measurements of 15 deep grove ball bearings of type LDK UER204 with a dynamic load rating of 12.82 kN. Measurements are performed on a test bench, where the bearings are subjected to accelerated degradation under different conditions. Three tests are performed with different loads and speeds, and five bearings are used in each test (see Tables 2 and 3). During the test, each bearing is instrumented with two piezoelectric accelerometers of type PCB 352C33 to collect vibration signals, mounted horizontally ($X$-axis) and vertically ($Y$-axis) and positioned at a 90 degree angle to each other. Vibration signals are captured at a sampling frequency of 25.6 kHz. The sampling frequency is 25.6 kHz, and 32,768 samples (i.e., 1.28 seconds of data) are recorded every 1 minute. We follow [63] and sample only from the $X$-axis; since the load is applied horizontally, the accelerometer positioned in this direction can capture more degradation information from the tested bearings. All bearings are run to failure under very high load to accelerate degradation. This significantly increases the risk of initiating bearing defects, but also the risk of local flash heating and subsequent uncontrolled damage to a bearing.
@@ 4.2 Feature extraction
We extract several features in the time domain from the raw bearing signal as important features for RUL prediction [59]. These features characterize the condition of mechanical systems based on sensor data. They include peak amplitude, energy, rise time, time, RMS, skewness, kurtosis, and crest factor, among others. Such features have been used before to train survival models with success [65], and have proven to be good predictors of the RUL [63]. In practice, the feature data come from an accelerometer placed on the ball bearings at an angle of 90 degrees. We record a total of 12 time-domain features for our empirical analyzes, which are summarized in Table 4.
@@ 4.3 Event detection algorithm
In this paper, we propose an event detection algorithm to estimate the time of failure before the end of life. We use this algorithm as an oracle to label the XJTU-SY dataset for model training and RUL prediction. The main motivation of our proposed algorithm stems from the idea that the recorded end of life, $L$, of the bearing constitutes an overestimation of the time when the bearing actually breaks. In other words, the bearing usually starts to fail or malfunction earlier than the time when the bearing becomes useless and the recording is stopped, and therefore, providing an inaccurate estimate of the breaking time is problematic because it can further bias the analysis.
In Algorithm (1), we present the pseudocode of our proposed event detection algorithm. In summary, the algorithm first performs the Hilbert transform and then applies the Fast Fourier Transform (FFT). This preliminary step is necessary to remove unwanted frequency components and eases the identification of the frequency bands of interest, as we explained in Section  3.1. The algorithm then systematically processes the signal obtained using a bandpass filter over the set of critical bands $CB=(cb_{i})_{i=1}^{5}$, as discussed in Section 3.1 and calculated based on the specific bearing properties of the XJTU-SY dataset in Table 3.
@@
where $w$ represents the window index, $\beta$ is an empirical deterioration parameter with units of $1/t$, $L$ is the end of life, and $\sigma_{KL}$ is the estimated standard variation of the measured differences of the KL divergence between windows. Finally, $\eta$ and $\lambda_{KL}$ are two parameters that are linked to the bearing’s natural deterioration and bear a physical interpretation.
In particular, the parameter $\eta$ reflects the sensitivity of the approach to changes at the beginning of the recording, which is linked to the estimated standard variation, $\sigma_{KL}$. We recommend setting this parameter to $\eta=5$ as it is sufficient to account for any initial variations until the bearing reaches a steady state. Note that this parameter may be set to higher values, yet this value is enough to avoid initial nonstationary perturbations. On the other hand, $\lambda_{KL}\sim 1$ controls how sensitive the algorithm is when it is close to the end of the bearing life. Unlike $\eta$, this additional parameter can change depending on the particular condition of the experiment, since each condition affects the bearing performance. For this experiment, we set $\lambda_{KL}=\{1.5,1.75,2\}$ for the high, medium and low operating conditions, respectively.
@@
Finally, the proposed KL-divergence metric provides better interpretability compared to deep learning architectures, such as autoencoders, because it has fewer parameters. Since our proposed method primarily relies on statistical and signal processing methods, there are no learned parameters (weights and biases) involved, making it more interpretable and parameter-efficient compared to a neural network. The primary adjustable parameters are $\eta$ and $\lambda_{KL}$, which influence detection sensitivity, but are not learned from the data.

### v3

@@ Abstract
Predicting the remaining useful life (RUL) of ball bearings is an active area of research, where novel machine learning techniques are continuously being applied to predict degradation trends and anticipate failures before they occur. However, few studies have explicitly addressed the challenge of handling censored data, where information about a specific event (e.g., mechanical failure) is incomplete or only partially observed. To address this issue, we introduce a novel and flexible method for early fault detection using Kullback-Leibler (KL) divergence and RUL estimation using survival analysis that naturally supports censored data. We demonstrate our approach in the XJTU-SY dataset using a 5-fold cross-validation strategy across three different operating conditions. When predicting the time to failure for bearings under the highest load (C1, 12.0 kN and 2100 RPM) with 25% random censoring, our approach achieves a mean absolute error (MAE) of 14.7 minutes (95% CI = 13.6-15.8) using a linear CoxPH model, and an MAE of 12.6 minutes (95% CI = 11.8-13.4) using a nonlinear Random Survival Forests model, compared to an MAE of 18.5 minutes (95% CI = 17.4-19.6) using a linear LASSO model that does not support censoring. Moreover, our approach achieves a mean cumulative relative accuracy (CRA) of 0.7586 over 5 bearings under the highest load, which improves over several state-of-the-art baselines. Our work highlights the importance of considering censored data as part of the model design when building predictive models for early fault detection and RUL estimation.
@@ 1 Introduction
However, despite numerous advances in model development and increases in predictive performances, one remaining challenge is how to handle censored data, where the event of interest is only partially observed – i.e., for some instances, we only have a lower bound on the actual time of the event. Censored or incomplete observations are a prevalent phenomenon given the rarity of failures or breakdowns in ball bearings [16], but so far, only few studies have attempted to model or address the consequences of censoring [69, 68, 23, 76] and the literature on censored data is limited. We can choose to ignore censored observations and analyze only the data where we know the true outcome, but this would lead to a loss of information and introduce an estimation bias in the model parameters [57]. If we want to accurately estimate when a bearing is likely to break down based on historical data and current sensor readings, we should include information from censored data samples in the model. Moreover, since bearing failure is generally a rare event, in the sense that the number of data samples where the bearing operates without problems far exceeds the number of samples with failure indications [16], it is important to make use of the data we have at our disposal, including censored and uncensored observations.
@@ 2.2 RUL prediction
[76] proposed a time-dependent survival neural network (TSNN), which estimates the risk of failure of a mechanical component by performing multiple classification steps in succession. The output of TSNN is an RUL-specific probability, which is monotonically decreasing. Their approach incorporates temporal dependencies between measurement points, and inherently supports censoring, but the method does not predict the time to failure, rather the probability of remaining failure-free by some time in the future. Additionally, the TSNN model requires a number of observational data points at prediction time. However, observational data require determining features beyond the observed time frame. As a result, time-variant predictions are difficult to make and survival bias may be introduced [80].
@@
[72] proposed a novel health indicator framework for automatic RUL prediction based on a multi-head attention architecture, combining information from damaged and healthy bearings to improve prediction accuracy. Their method achieves a good prediction error compared to other works, but the estimated RUL is not monotonically decreasing, and the authors make the assumption of a linear correlation between the RUL and a bearing’s lifetime. Assuming a linear RUL oversimplifies the bearing degradation process, as the relationship between the RUL and the number of life cycles is often nonlinear [30, Ch. 2]. However, bearings may not degrade at a constant speed, as their health can depend on various factors, e.g., operating condition, the type of maintenance performed, or other external factors [78].
@@
[39] proposed a hybrid deep learning approach for RUL prediction that utilizes non-full lifecycle (i.e., censored) data. The authors combine self-supervised and supervised learning to address data scarcity and noise challenges in RUL prediction tasks. Using a Contrastive Predictive Coding (CPC) encoder and a Transformer decoder, the model leverages non-full lifecycle data for pretraining and finetuning with full lifecycle data. Their experiments show good performance in a separate dataset in terms of the root mean squared error (RMSE). Although censored data are used in the pretraining phase, the prediction task in this case is simply predicting the next feature value $x_{t+1}$ given $x_{t}$, not the time-to-event. In other words, the censored data is not used to predict the probability of future event, but rather to improved the learned feature presentation. Moreover, the authors never evaluate their method with censoring in the testing phase and the proposed RMSE assumes that all samples have observed failure times and the predicted RUL is not monotonically decreasing.
@@ 3.2 Defect evolution and failure monitoring
The early stages of a bearing defect can be effectively detected from weak vibration signals through the enveloped Fourier transform [50], as this method can detect amplitude-modulated signals. By selecting an impulsive frequency band, often containing one or more of the component’s eigenfrequencies and applying the Hilbert transform, the low-frequency envelope or carrier frequency can be found [49, Ch. 3]. The carrier frequency corresponds to one of the four characteristic bearing frequencies that are bound to the bearing geometry, the number of rolling elements, and the rotation speed.
@@ 3.3 Elements of survival analysis
We define a survival problem by a sequence of observations represented as triplets ($\bm{x}_{i}$, $t_{i}$, $\delta_{i}$), where $\bm{x}_{i}\in\mathbb{R}^{d}$ is a vector with $d$ features, $t_{i}$ denotes the time-to-event or censoring, and $\delta_{i}$ is a binary event indicator. Moreover, let $c_{i}$ denote the censoring time and $e_{i}$ denote the event time for the $i$th observation, thus $t_{i}=e_{i}$ if $\delta_{i}=1$ or $t_{i}=c_{i}$ if $\delta_{i}=0$. We consider the survival time as discrete and limit the time horizon to a finite duration, denoted $(0,...,T_{max})$, where $T_{max}$ represents a predefined maximum time point (e.g., 1 year). A survival model then estimates the probability that the event of interest occurs at time $T$ later than $t$, that is, the survival probability $S\left(t\right)=\text{P}\left(T>t\right)=1-\text{P}\left(t\leq T\right)$. To estimate the survival probability, we can use the hazard function:
@@
which represents the failure rate at an instant after time $t$, assuming survival beyond that time [19, Ch. 11]. The hazard function is related to the survival function through $h\left(t\right)=f\left(t\right)/S\left(t\right)$, where $f\left(t\right)$ is the probability density associated with $T$. Formally, $f\left(t\right):=\lim_{\Delta t\rightarrow 0}\text{P}\left(t<T\leq t+\Delta t\right)/\Delta t$, that is, the instantaneous failure rate at time $t$. Under this definition, the function $h\left(t\right)$ is the probability density of $T$ conditioned on $T>t$, where the higher the value of $h\left(t\right)$, the higher the probability of failure. The functions $S\left(t\right)$, $h\left(t\right)$, and $f\left(t\right)$, are different but related ways to describe the probability distribution of $T$.
@@ 3.4 Elements of uncertainty estimation
Many machine learning methods for RUL prediction are indeed nonprobabilistic; they can only provide a point estimate of the RUL and disregard the inherent uncertainty associated with the prediction. The Bayesian framework offers a theoretical approach to model uncertainty when estimating the model parameters (epistemic) and when making the actual prediction (aleatoric) [41]. This enables us to assess confidence in the predictions, which further encourages the adoption of the model as a decision support tool. [44] introduced Bayesian Neural Networks (BNNs) for bearing health prognostics with uncertainty estimation to predict the RUL of turbofan engines. In a BNN, the weights are treated as a set of random variables $\bm{\theta}\sim\text{P}\left(\bm{\theta}\right)$. In this context, let $\mathcal{D}$ denote the data, $\text{P}\left(\mathcal{D}|\bm{\theta}\right)$ the likelihood and $\text{P}\left(\bm{\theta}\right)$ the prior distribution over the parameters of interest $\bm{\theta}$. Bayesian inference computes the posterior distribution $\text{P}\left(\bm{\theta}|\mathcal{D}\right)=\text{P}\left(\bm{\theta}\right)\text{P}\left(\mathcal{D}|\bm{\theta}\right)/\text{P}\left(\bm{\theta}\right)$, and then obtains the predictive distribution of unobserved data conditioned on $\mathcal{D}$ by marginalizing over the parameter space $\Theta$, i.e.$\text{P}\left(\bm{x}_{\text{new}}|\mathcal{D}\right)=\int_{\Theta}\text{P}\left(\bm{x}_{\text{new}}|\mathcal{D},\bm{\theta}\right)\text{P}\left(\bm{\theta}|\mathcal{D}\right)\text{d}\bm{\theta}$. However, the posterior cannot be directly computed due to $\text{P}\left(\mathcal{D}\right)$ being intractable, and sampling methods do not scale well in high dimensions. To this end, variational inference (VI) has been proposed to approximate the posterior with a tractable parametric distribution, $q_{\bm{\zeta}}\left(\bm{\theta}\right)$. This $q_{\bm{\zeta}}\left(\bm{\theta}\right)$ is chosen within a class of tractable parametric distributions, $\mathcal{Q}$, by minimizing the KL divergence from $q_{\bm{\zeta}}\left(\bm{\theta}\right)$ to $\text{P}\left(\bm{\theta}|\mathcal{D}\right)$. In many modern neural networks, Monte Carlo Dropout (MCD) or another timely technique is used to approximate the VI solution [18][36] proposed BNNs as a tool for uncertainty estimation in deep survival networks using MCD and showed a performance advantage in adopting the Bayesian framework, especially in small datasets, over nonprobabilistic models. The intrinsic probabilistic dimension of BNNs naturally allows us to make uncertainty estimates, estimate predictive errors, and plot credible intervals [47]. This, in turn, can justify the level of trust and confidence that practitioners have in the prediction. Nonprobabilistic models cannot give these advantages.
@@
Figure 2: Outline of the proposed method. Historical data for 5 ball bearings form a bearing dataset $\mathcal{D}$ (see Section 4.1). We first perform feature extraction of all the bearings in the time-domain (see Section 4.2). Then, to train a survival model for RUL prediction, we designate training and test bearings and perform event detection in the frequency-domain of the training bearings (see Section 4.3). Afterwards, we convert the temporal training dataset to a supervised learning dataset by computing a rolling average (see Section 4.4. We then apply random, independent censoring using the proposed censoring algorithm (see Section 4.5), so that the event of interest is only observed for a portion of the instances. This processed dataset $\tilde{\mathcal{D}}$ thus contains $N$ instances (rows) with several time-domain features (e.g., entropy, kurtosis), and the time-to-event (filled dot) or censoring (hollow dot). We use this dataset (features and event information) to train a survival model $\mathcal{M}$ (see Section 4.6) that can estimate the individual survival distribution (ISD) of the failure event, for a new test bearing $\bm{x}_{i}$, denoted as $\hat{s}_{i}$. This ISD give the probability that failure (i.e., the onset of degradation) occurs after $t$ minutes post startup, for all $t>0$. It can also be used to estimate the time-to-event for this $\bm{x}_{i}$, for example, when the survival (event) curve intersects the dashed horizontal line at 50%, which is referred to as the predicted median survival time.
@@ 4.1 The dataset
We use the publicly-available Xi’an Jiaotong University (XJTU-SY) [64] dataset to evaluate the performance of our proposed method. The dataset contains measurements of 15 deep grove ball bearings of type LDK UER204 with a dynamic load rating of 12.82 kN. Measurements are performed on a test bench, where bearings are subjected to accelerated degradation under different conditions. Three tests are performed with different loads and speeds, and five bearings are used in each test (see Tables 2 and 3). During the test, each bearing is instrumented with two piezoelectric accelerometers of type PCB 352C33 to collect vibration signals, mounted horizontally ($X$-axis) and vertically ($Y$-axis) and positioned at a 90 degree angle to each other. Vibration signals are captured at a sampling frequency of 25.6 kHz. The sampling frequency is 25.6 kHz, and 32,768 samples (i.e., 1.28 seconds of data) are recorded every 1 minute. We follow [64] and sample only from the $X$-axis; since the load is applied horizontally, the accelerometer positioned in this direction can capture more degradation information from the tested bearings. All bearings are run to failure under very high load to accelerate degradation. This significantly increases the risk of initiating bearing defects, but also the risk of local flash heating and subsequent uncontrolled damage to a bearing.
@@ 4.2 Feature extraction
We extract several features in the time domain from the raw bearing signal as important features for RUL prediction [60]. These features characterize the condition of mechanical systems based on sensor data. They include peak amplitude, energy, rise time, time, RMS, skewness, kurtosis, and crest factor, among others. Such features have been used before to train survival models with success [66], and have been shown to be good predictors of the RUL [64]. In practice, the feature data come from an accelerometer placed on the ball bearings at an angle of 90 degrees. We record a total of 12 time-domain features for our empirical analyzes, which are summarized in Table 4.
@@ 4.3 Event detection algorithm
In this paper, we propose an event detection algorithm to estimate the time of failure before the end of life. We use this algorithm as an oracle to label the XJTU-SY dataset for model training and RUL prediction. The main motivation for our method comes from the idea that the recorded end of life, $L$, of the bearing constitutes an overestimation of the time when the bearing actually breaks. In other words, the bearing usually starts to fail or malfunction earlier than the time when the bearing becomes useless and the recording is stopped, and therefore, providing an inaccurate estimate of the breaking time is problematic because it can further bias the analysis.
In Algorithm (1), we present the pseudocode of our proposed event detection algorithm. In summary, the algorithm first performs the Hilbert transform and then applies the Fast Fourier Transform (FFT). This preliminary step is necessary to remove unwanted frequency components and eases the identification of the frequency bands of interest, as we explained in Section  3.1. The algorithm then systematically processes the signal obtained using a band-pass filter over the set of critical bands $CB=(cb_{i})_{i=1}^{5}$, as discussed in Section 3.1 and calculated based on the specific bearing properties of the XJTU-SY data set in Table 3.
@@
where $w$ represents the window index, $\beta$ is an empirical deterioration parameter with units of $1/t$, $L$ is the end of life, and $\sigma_{KL}$ is the estimated standard variation of the measured differences of the KL divergence between windows. Finally, $\eta$ and $\lambda_{KL}$ are two parameters that are linked to the natural deterioration of the bearing and bear a physical interpretation.
In particular, the parameter $\eta$ reflects the sensitivity of the approach to changes at the beginning of the recording, which is linked to the estimated standard variation, $\sigma_{KL}$. We recommend setting this parameter to $\eta=5$ as it is sufficient to account for any initial variations until the bearing reaches a steady state. Note that this parameter may be set to higher values, yet this value is enough to avoid initial nonstationary perturbations. On the other hand, $\lambda_{KL}\sim 1$ controls how sensitive the algorithm is when it is close to the end of the bearing life. Unlike $\eta$, this additional parameter can change depending on the particular condition of the experiment, since each condition affects the bearing performance. For this experiment, we set $\lambda_{KL}=\{1.5,1.75,2\}$ for the high, medium, and low operating conditions, respectively.
@@
Finally, the proposed KL-divergence metric provides better interpretability compared to deep learning architectures, such as autoencoders, because it has fewer parameters. Since our proposed method relies primarily on statistical and signal processing methods, there are no learned parameters (weights and biases) involved, making it more interpretable and parameter-efficient compared to a neural network. The primary adjustable parameters are $\eta$ and $\lambda_{KL}$, which influence the detection sensitivity, but are not learned from the data.